OpenStudy (anonymous):

Find the general solution of the differential equation 3x dy/dx=2y and then find the value of C if when x=8, y=2. finally write the particular solution of the equation for the indicated condition.

OpenStudy (anonymous):

y(x)=cx^(2/3) c=1/2

OpenStudy (anonymous):

how would you solve the integral of 3x/dx? because i got that equal to the integral of 2ydy